# Distributed Solving of Linear Quadratic Optimal Controller with Terminal State Constraint

### v3

@@ 6 Conclusions
This paper studied the LQ optimal control problem for continuous-time systems with terminal state constraints. A distributed algorithm was proposed to solve the optimal controller where only partial information is accessible to the agent, consisting of the distributed iteration of the Riccati equation, the optimal Lagrange multiplier, and the optimal state. Numerical examples demonstrated the effectiveness of the proposed algorithm.
@@
$\displaystyle\lim_{k\rightarrow\infty}\|Z_{i,k}^{1}(t)-Z(t)\|$ $=$ $0,$ (88)
$\displaystyle\lim_{k\rightarrow\infty}\|V_{i,k}^{1}(t)-V(t)\|$ $=$ $0$ (89)
@@
$\displaystyle\lim_{k\rightarrow\infty}\|V_{i,k}^{1}(t)-V_{j,k}^{1}(t)\|$ $=$ $0$ (91)
@@
$\displaystyle\lim_{k\rightarrow\infty}\|\frac{1}{N}\sum_{i=1}^{N}V_{i,k}^{1}(t)-V^{1}(t)\|$ $=$ $0$ (93)
@@
$\bar{\mathcal{Z}}_{k}^{1}(t)$ $=$ $[(I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{Z}}_{k-1}^{1}(t)$
@@
$\bar{\mathcal{V}}_{k}^{1}(t)$ $=$ $[(I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{V}}_{k-1}^{1}(t)$
@@
×$\times[(I_{N}-\mathcal{B})\otimes I_{n}]\Pi_{\tau-1}^{1}(t),$
@@
×$\times[(I_{N}-\mathcal{B})\otimes I_{n}]\Theta_{\tau-1}^{1}(t).$
@@
×$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Pi_{\tau-1}^{1}(t)\|,$
@@
×$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|.$
@@
×$\quad\quad\quad\quad\quad\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Pi_{\tau-1}^{1}(t)\|=0,$ (98)
@@
×$\quad\quad\quad\quad\quad\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|=0.$ (99)
@@
By applying the facts that $10<1-\alpha_{k}<$ 1, and $\lim_{k\to\infty}\alpha_{k}=0$, it yields that

### v4

@@ VII Conclusions
This paper studied the LQ optimal control problem for continuous-time systems with terminal state constraints. A distributed algorithm was proposed to solve the optimal controller where only partial information is accessible to the agent, consisting of the distributed iteration of the Riccati equation, the optimal Lagrange multiplier, and the optimal state. As a significant application, the proposed distributed approach is extended to address the optimal consensus control problem. Numerical examples demonstrated the effectiveness of the proposed algorithm, where the performance index under the proposed distributed controller is smaller than that under the commonly used consensus control.
@@
$\displaystyle\lim_{k\rightarrow\infty}\|Z_{i,k}^{1}(t)-Z^{1}(t)\|$ $=$ $0,$ (87)
$\displaystyle\lim_{k\rightarrow\infty}\|V_{i,k}^{1}(t)-V^{1}(t)\|$ $=$ $0.$ (88)
@@
$\displaystyle\lim_{k\rightarrow\infty}\|V_{i,k}^{1}(t)-V_{j,k}^{1}(t)\|$ $=$ $0.$ (90)
@@
$\displaystyle\lim_{k\rightarrow\infty}\|\frac{1}{N}\sum_{i=1}^{N}V_{i,k}^{1}(t)-V^{1}(t)\|$ $=$ $0.$ (92)
@@
$\bar{\mathcal{Z}}_{k}^{1}(t)$ $=$ $[(I-\alpha_{k}I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{Z}}_{k-1}^{1}(t)$
@@
$\bar{\mathcal{V}}_{k}^{1}(t)$ $=$ $[(I-\alpha_{k}I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{V}}_{k-1}^{1}(t)$
@@
$\times[(I_{N}-\mathcal{B})\otimes I_{n}]\Pi_{\tau-1}^{1}(t),$
@@
$\times[(I_{N}-\mathcal{B})\otimes I_{n}]\Theta_{\tau-1}^{1}(t).$
@@
$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Pi_{\tau-1}^{1}(t)\|,$
@@
$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|.$
@@
$\quad\quad\quad\quad\quad\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Pi_{\tau-1}^{1}(t)\|=0,$ (97)
@@
$\quad\quad\quad\quad\quad\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|=0.$ (98)
@@
By applying the facts that $0<1-\alpha_{k}<1$, and $\lim_{k\to\infty}\alpha_{k}=0$, it yields that